### v1

# MOHPER: Multi-objective Hyperparameter Optimization Framework for E-commerce Retrieval System

Abstract.
E-commerce search optimization has evolved to include a wider range of metrics that reflect user engagement and business objectives. Modern search frameworks now incorporate advanced quality features, such as sales counts and document-query relevance, to better align search results with these goals. Traditional methods typically focus on click-through rate (CTR) as a measure of engagement or relevance, but this can miss true purchase intent, creating a gap between user interest and actual conversions. Joint training with the click-through conversion rate (CTCVR) has become essential for understanding buying behavior, although its sparsity poses challenges for reliable optimization. This study presents MOHPER, a Multi-Objective Hyperparameter Optimization framework for E-commerce Retrieval systems. Utilizing Bayesian optimization and sampling, it jointly optimizes both CTR, CTCVR, and relevant objectives, focusing on engagement and conversion of the users. In addition, to improve the selection of the best configuration from multi-objective optimization, we suggest advanced methods for hyperparameter selection, including a meta-configuration voting strategy and a cumulative training approach that leverages prior optimal configurations, to improve speeds of training and efficiency. Currently deployed in a live setting, our proposed framework substantiates its practical efficacy in achieving a balanced optimization that aligns with both user satisfaction and revenue goals.
Retrieval System, Multi-objective Hyperparameter Optimization
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Retrieval models and rankingccs: Computing methodologies Learning paradigmsccs: Computing methodologies Rankingccs: Applied computing Online shopping
1. Introduction
With the growth of online shopping in modern society, e-commerce platforms now account for a vast number of user queries daily. The product search engines within these platforms play a pivotal role, acting as arbitrators that match consumer intent with product listings. Beyond simply interpreting and retrieving relevant items, e-commerce search engines serve as the primary conduit between digital vendors and consumers, shaping the user experience and influencing engaging and purchasing behavior across online markets. To improve both the engagement of consumers and the profitability of vendors, e-commerce search optimization has rapidly progressed from simple keyword matching to sophisticated, multi-feature frameworks that integrate relevance, user engagement, and business-oriented metrics (Wu et al., 2018; Zheng et al., 2022; Karmaker Santu et al., 2017).
Initially, search systems relied heavily on keyword-matching algorithms such as TF-IDF, and BM25, which were limited to capturing direct lexical overlap between search keywords and product listings (Ramos et al., 2003; Robertson et al., 2009). However, the recent breakthrough of large language models (LLMs) or neural multi-modal retrieval models makes systems process dense retrievals, by measuring semantic similarities of query-document vectors, as introduced by several studies (Arevalillo-Herráez et al., 2013; Li et al., 2023a; Bhagdev et al., 2008). Although most previous studies focus on relevance scores of the hybrid retrieval using both BM25 and dense retrieval, these similarity scores do not imply the preferences, interests, or engagements of users, which are crucial for e-commerce, generally available by interpreting user interaction logs. To leverage user activity log data, such as selling, clicks, etc. from actual e-commerce service, e-commerce search optimization has evolved markedly, moving from traditional keyword-based retrieval methods to advanced frameworks incorporating relevance, engagement, and an array of business-critical metrics.
Figure 1. Overall architecture diagram of the proposed framework, MOHPER.
[…]
architecture
Historically, search optimization has centered on maximizing click-through rate (CTR) as an engagement metric, but CTR alone is often insufficient to predict actual conversions, a critical focus in e-commerce. To address this, click-through conversion rate (CTCVR) has become essential for capturing purchase intent. However, training solely on CTCVR presents challenges due to sample selection bias and data sparsity, where training data derived from clicks does not generalize well to the broader space of all impressions, and conversion data remains limited, risking overfitting to a small subset of items (Zhang et al., 2022; Ma et al., 2018). Traditional retrieval or ranking methods, including BM25 (Robertson et al., 2009), dense vector retrievals (Wang et al., 2021; Li et al., 2023a, b), and learning-to-rank (LTR) (Burges et al., 2005), typically optimize for single objectives, which can result in misalignment between relevance, user engagement, and conversion. This single-objective focus limits adaptation to rapidly changing e-commerce trends and degrades performance in online environments (Karmaker Santu et al., 2017).
In response, multi-objective strategies are gaining traction in e-commerce, aiming to optimize both CTR and CTCVR to align engagement with purchase intent. This study presents MOHPER, a multi-objective hyperparameter optimization (MOHPO) framework designed for e-commerce search. Using Bayesian optimization as (Li and Kanoulas, 2018) and advanced sampling techniques, MOHPER jointly optimizes CTR and CTCVR, offering a balanced solution to enhance both engagement and conversions. We systematically evaluate sampling strategies within MOHPER to ensure robust multi-metric optimization, addressing the demand for adaptive, multi-objective solutions in e-commerce. The main contributions of this paper include:
  1. i.
    introduction of a novel hyperparameter optimization (HPO) framework specifically tailored for e-commerce search services, focusing on objectives such as CTR and CTCVR essential indicators of user engagement and purchase intent. This is the first application of a multi-objective HPO framework in e-commerce retrieval and ranking tasks.
  2. ii.
    suggestion for advanced methodologies for hyperparameter selection, including a meta-configuration voting strategy and a cumulative training approach that leverages prior optimal configurations.
  3. iii.
    thorough testing of the MOHPER framework using various samplers and search scenarios, validated through extensive A/B testing with a significant dataset from OHouse, one of South Korea’s leading furniture e-commerce platforms.
The notable performance improvements indicate that models optimized via MOHPER have been successfully adopted as foundational models on the OHouse platform, showcasing the practical effectiveness of our approach in real-world e-commerce applications.
2. Proposed Framework: MOHPER
In this section, we present the proposed MOHPER framework. We begin by defining the multi-objective problem within retrieval systems and provide an overview of the HPO algorithm. Subsequently, we identify the relevant objectives for joint optimization in the e-commerce setting. To enhance performance during the meta-evaluation stage and to select appropriate configurations in the multi-objective optimization context, we introduce two methods: meta-configuration voting and cumulative hyperparameter stacking. The proposed methods are illustrated in Figure 1.
2.1. Multi-objective HPO in Retrieval System
For MOHPO, the objective is to identify a set of HPs that optimizes the performance across multiple objectives. These objectives typically conflict, meaning improving one objective may negatively impact another. The goal of HPO is to find an optimal configuration of HPs that balances these objectives. Similarly, in the context of a retrieval system, MOHPO aims to improve the system’s ability to return relevant results for a given set of queries. The objectives generally focus on relevance, ranking quality, and overall system performance. Some common performance metrics for retrieval systems can be either normalized discounted cumulative gain (nDCG), mean average precision (mAP), precision, or recall (Järvelin and Kekäläinen, 2002; Zhu, 2004) while optimizing different labels indicating relevance level, user engagement, or freshness of items, etc.
Figure 2. Illustration of transform function to modulate query request for a search engine as the change of config hyperparameter.
[…]
ModelInput
Formally, the goal is to identify a set of HPs ($h\in\mathcal{H}$) that maximizes the performance of a retrieval engine across multiple objectives $O=\{O_{1},O_{2},\dots,O_{M}\}$ simultaneously. Given an observation dataset $\mathcal{D}_{0}=\emptyset$ and a set of ground truth labels $Y=\{y_{1,(\cdot)},y_{2,(\cdot)},\dots,y_{m,(\cdot)}\}$ for each objective, the optimization seeks to sample and evaluate HP configurations to approximate optimal performance iteratively. In other words, the search system is evaluated by selecting an HP configuration $h_{t}$ from the search space $\mathcal{H}$ at each trial $t=1,2,\dots,N$. During optimization progress, various sampler methods, such as random search, grid search, or Bayesian optimization (e.g. tree-structured Parzen estimator (TPE) or Gaussian process (GP), etc.), can be employed to sample HP configurations. At each trial, $t$, an HP configuration $h_{t}$ is sampled, and the retrieval system is evaluated on a set of queries. The configuration $h_{t}$ influences the retrieval process through transform functions that adapt the retrieval engine’s behavior based on the sampled HPs. This could involve altering ranking strategies, modifying relevance weights, or adjusting search heuristics (see Figure 2). Specifically, the transform function can be expressed as $T(h_{t},q)$, which represents the transformed retrieval process based on the HP configuration $h_{t}$ applied to query $q$. Depending on the sampled HPs, this transformation could modify how items are ranked or filtered for search engines. For each query $q$, the system produces predictions (ranking of items), and the metrics of relevance/quality for these items are evaluated using objectives, respectively. The metric for objective $m$ can be computed based on the predicted rankings $\hat{y}_{h_{t},q,(\cdot)}$ and compared to the ground truth labels $y_{m,q,(\cdot)}$, formulated as $z_{t,m,q}=O_{k}(y_{m,q,(\cdot)},\hat{y}_{h_{t},q,(\cdot)})$. The performance evaluations across all queries are then aggregated to produce a performance metric for each objective. The aggregated metrics are used to update the optimization process, guiding the search for better hyperparameter configurations. This process continues iteratively until convergence, where the goal is to select a hyperparameter configuration that optimizes the trade-off between the multiple objectives. The overall summarized algorithm with TPE sampler is summarized in Algorithm 1.
Algorithm 1 MOHPER Training with the TPE Sampler Example
1:Initialize HPO study with sampler $S$ (e.g., TPE, GP, …)
2:Define objectives $O=\{O_{1},O_{2},\dots,O_{M}\}$, with corresponding ground truth labels $Y=\{y_{1,(\cdot)},y_{2,(\cdot)},\dots,y_{M,(\cdot)}\}$
3:If it is in cumulative training, load previous top observations into datasets $\mathcal{D}_{0}=\{h_{prev\ A},z_{prev\ A,1},z_{prev\ A,2},\cdots\}$ in cumulative training, else initialize observation dataset $\mathcal{D}_{0}=\emptyset$
4:for each trial $t=1,2,\dots,N$ do
5:     Sample hyperparameters (HPs):
$h_{t}\sim S(\mathcal{H},\mathcal{D}_{t-1},\text{params})$
6:     for each query $q=q_{1},q_{2},\dots,q_{L}$ do
7:         Retrieve Items from Search Engine ($R$) with Transform Function ($T$):
$\hat{y}_{h_{t},q,(\cdot)}=R(T(h_{t},q,(\cdot)))$
8:         Evaluate objectives:
$z_{t,m,q}=O_{m}(y_{m,q,(\cdot)},\hat{y}_{h_{t},q,(\cdot)}),\quad\forall O_{m}\in O$
9:     end for
10:     Aggregate evaluated objectives (e.g., average):
$z_{t,m}=Agg(z_{t,m,q}),\quad\forall O_{m}\in O$
11:     Store results $(h_{t},z_{t,1},z_{t,2},\dots,z_{t,M})$
12:     Update TPE sampler state based on evaluations:
13:     Define threshold $\gamma_{t,m}$ on each objective, for $v$ quantile:
$\gamma_{t,m}=\text{quantile}(z_{t,m},v),\quad\forall O_{m}\in O$
14:     Update good and bad HP distributions using $\gamma_{t,m}$:
$l(\mathbf{h})=p(\mathbf{h}\mid z_{t,m}\leq\gamma_{t,m}),\quad g(\mathbf{h})=p(\mathbf{h}\mid z_{t,m}>\gamma_{t,m})$
15:     Calculate expected improvement (EI) for candidate HPs and select HPs that maximize EI:
$\text{EI}(\mathbf{h})=\prod_{m=1}^{M}\frac{l(\mathbf{h})}{g(\mathbf{h})},\quad h_{t}=\arg\max_{\mathbf{h}\in\mathcal{H}}\text{EI}(\mathbf{h})$
16:     Update dataset if improvement is achieved: If $\text{EI}(h_{t})>0$, then update the observation dataset:
$\mathcal{D}_{t}=\mathcal{D}_{t-1}\cup\{(h_{t},z_{t,1},z_{t,2},\dots,z_{t,M})\}$
17:end for
2.2. Joint Optimization in E-Commerce Context
Figure 3. Illustration of the e-commerce conversion funnel starting from item impressions, highlighting issues of sample selection bias and data sparsity in user interaction logs
[…]
Illustration of sample selection bias and data sparsity issue
To enhance user engagement and sales in e-commerce while addressing sample selection bias and data sparsity, it is essential to mine user interaction logs, including clicks and conversions per impression. Although MOHPO is designed to optimize any metrics of objectives as other black-box HPO algorithms do, in this study, we track the retrieval performance through the following objectives:
  1. (1)
    Click-Through Conversion Rate (CTCVR): =#$=\frac{\#purchase}{\#imp}$
  2. (2)
    Click-Through Rate (CTR) =#$=\frac{\#click}{\#imp}$
  3. (3)
    Click-Through Cart-Adding Rate (CTCAR): =#$=\frac{\#cart}{\#imp}$
Here, #$\#imp$ represents the number of item impressions in retrievals. CTCVR is the primary objective, with CTR, and CTCAR serving as auxiliary objectives to mitigate selection bias and sparsity (Figure 3) (Ma et al., 2018; Karmaker Santu et al., 2017). Specifically, For any query $q$ and trial $t$, performance metrics $z_{t,q,(\cdot)}$ (e.g. precision@K, nDCG@K, mAP@K, and recall@K) are computed when ground truth labels $y_{q,(\cdot)}$ are available and the top-K items are retrieved ($\hat{y}_{h_{t},q,(\cdot)}$). Labels for these items are derived from objective-specific statistics (e.g. #$\#imp_{q,i}$, #$\#click_{q,i}$, #$\#purchase_{q,i}$), with thresholding for positive/negative scoring at the corresponding objective. Also, while using naive statistical rates is available, filtering based on impression counts and Bayesian smoothing for rare events can further refine these scores (Wang et al., 2011). Following metric calculation $z_{t,m,q}$, metrics are aggregated across queries, and MOHPO proceeds with multi-objective tuning to enhance engagement, conversion efficiency, and purchase intent, advancing retrieval performance in e-commerce.
2.3. Hyperparameter Extraction Stage
Upon completing HPO training, the framework selects a diverse set of candidate hyperparameter configurations based on their performance across multiple objectives, such as training targets or a weighted metric sum (see the middle stage of Figure 1). These top configurations are then stored for further meta-level evaluation to estimate online performance and for use in subsequent cumulative training (see subsection 2.5).
2.4. Meta Stage Evaluation and HP selection
Selecting hyperparameter configurations based solely on training metrics can lead to overfitting, as HPO is often biased toward training data and may not generalize well to production due to sample selection bias and data sparsity issues. To address this, the meta-level evaluation assesses top-performing HP configurations on unseen queries and future periods using time-based and query-wise data splits, ensuring no overlap with training data. Each HP is evaluated for robustness across multiple metrics, and we employ a meta-configuration voting strategy (Chan and Stolfo, 1995) to prioritize HPs that demonstrate consistent, balanced performance across objectives. This approach synthesizes the strengths of various configurations, ensuring the selection of parameters that generalize effectively.
2.5. Cumulative Training
While the baseline algorithm (Sections 3.1–3.4) can operate independently, incorporating prior optimal hyperparameters into each successive MOHPO iteration enhances both performance and stability. Our cumulative training approach initializes each new training cycle with a dataset $\mathcal{D}_{0}=\{h_{prev\ A},z_{prev\ A,1},z_{prev\ A,2},\cdots\}$ in cumulative training, instead of initializing observation dataset $\mathcal{D}_{0}=\emptyset$. This setup allows the sampler to concentrate on unexplored regions, minimizing redundant samples and promoting exploration within the configuration space.
This cumulative method effectively prioritizes configurations with consistently strong performance across objectives, reducing computational redundancy and adapting dynamically to changes in user behavior and market trends. By establishing an adaptive feedback mechanism, cumulative training fortifies the framework’s robustness in evolving e-commerce environments.
3. Experiments
We conduct extensive experiments to evaluate the effectiveness and the performance of the MOHPO framework and its components and the following research questions (RQs) are answered:
  1. RQ 1.
    Multi-Objective Balancing Across Objectives and Samplers: How effectively does MOHPO balance multi-objective metrics (e.g., CTR, CTCVR, CTCAR) under varying objective configurations and HPO sampling strategies?
  2. RQ 2.
    Hybrid Search Optimization with MOHPO: How well does MOHPO integrate hybrid search mechanisms to optimize for both relevance and multi-objective goals, enhancing user engagement and conversion?
  3. RQ 3.
    Effectiveness of Meta-Configuration Voting and Cumulative Training: Do meta-configuration voting and cumulative training improve MOHPO’s generalizability on unseen queries and adaptability to temporal shifts? Are these strategies critical for maintaining high performance in dynamic e-commerce environments?
  4. RQ 4.
    Online Deployment and Practical Feasibility: Can MOHPO be effectively deployed in real-world online services while ensuring computational efficiency and demonstrating tangible improvements in retrieval/ranking performance?
3.1. Datasets and Experimental Setups
Table 1. The basic statistics of datasets using train/test.
Dataset # Queries # Impressions # Clicks # Conversions
SERP Sampled Train 1,000 367,544,388 12,021,645 1,214,896
SERP Sampled Test 5,000 377,367,483 11,978,753 1,230,871
[…]
CPLP Test 1,255 106,484,599 2,588,947 228,671
  • SERP refers to the search engine result page, employing keywords as queries, and CPLP means the product list page, utilizing category IDs for filtering products. The dataset comprises a subset of production data for baseline comparisons. For online training, the other queries, including tail queries, are integrated into the offline dataset to ensure comprehensive training/evaluation (5,000 for training and 30,000 for testing).
3.1.1. Datasets and Split Settings
For offline evaluation, we utilize user interaction logs from the OHouse [^1] https://ohou.se/ platform, capturing diverse engagement events such as clicks, purchases, and cart additions across a broad spectrum of queries. In this study, we draw a subset of interaction logs from users in the Republic of Korea spanning from June 16 to September 15, 2024, for training data. The subsequent month’s logs (September 16 to October 15, 2024) are employed for meta-stage evaluation. Offline evaluation is conducted across two scenarios: (1) a keyword-search scenario, where queries are based on textual input, and (2) a retrieval scenario, where queries are formulated through product category IDs, representing groups such as furniture, electronics, and kitchen goods. The statistics of the training and evaluation datasets are summarized in Table 1.
3.1.2. Retrieval Engine and Indexing Setting
For our experiments, we employ Elasticsearch (ES) [^2] https://github.com/elastic/elasticsearch, specifically version 8.14 for offline training and evaluation tasks and version 8.8 for online experiments. Built upon the Apache Lucene library (McCandless et al., 2010; Białecki et al., 2012), ES is a real-time, scalable, distributed, and multi-tenant search engine supporting both textual and vector search capabilities, widely adopted by several e-commerce websites. Textual features are indexed via inverted indices (Salton et al., 1983), while vector features leverage hierarchical navigable small-world (HNSW) indexing (Malkov and Yashunin, 2018). For experimental consistency, we maintain separate ES instances for offline evaluation and production use, with partial item indices from production duplicated for experimental use. These indices are configured on 3 primary shards with 1 replica each, yielding a total of 6 shards.
Given the multiple retrieval operations required per training trial in HPO, we use the multi-search API [^3] https://www.elastic.co/guide/en/elasticsearch/reference/current/search-multi-search.html to execute concurrent searches within a single request, similar to a mini-batch bulk request approach. This concurrent processing is essential for the HPO process, where search results are repeatedly requested to generate predictions based on various hyperparameter configurations, significantly improving training efficiency through parallelization.
3.1.3. Hyperparamter Optimization Framework
While various hyperparameter optimization (HPO) frameworks can be applied to MOHPER, we utilize Optuna version 4.0 (Akiba et al., 2019), enabling the evaluation of diverse samplers and configuration settings. Optuna’s flexibility and adaptability make it well-suited for the rigorous testing of multiple HPO strategies within MOHPER. To address the need for concurrent and parallel processing across multiple training trials, we integrate RayTune [^4] https://docs.ray.io/en/latest/tune/index.html, built on the Ray framework [^5] https://github.com/ray-project/ray, which is designed to support scalable, distributed computing within Python [^6] https://www.python.org/ applications (Liaw et al., 2018). This integration supports high-throughput, parallelized experimentation, enhancing the efficiency of testing various sampler configurations within MOHPER.
Table 2. Performance comparison of various models and optimization objectives on SERP test data. For sampling-based HPO (excluding grid search), 100 samples are limitedly used by default to show sample efficiencies unless otherwise specified.
Method CTR nDCG@20 CTR nDCG@100 CTR prec@100 CTR avg metrics CTCVR nDCG@20 CTCVR nDCG@100 CTCVR prec@100 CTCVR avg metrics Overall avg metrics
BM25 0.2273 0.2630 0.2724 0.2542 0.0779 0.1470 0.3977 0.2076 0.2309
ANN (KoBERT) 0.5295 0.3592 0.2836 0.3908 0.2799 0.3032 0.4027 0.3286 0.3597
ANN (Jina-clip-v2, query to item text) 0.4071 0.3206 0.2792 0.3356 0.2217 0.2643 0.4349 0.3070 0.3213
ANN (Jina-clip-v2, query to item image) 0.4539 0.3370 0.2818 0.3576 0.2465 0.2825 0.4246 0.3179 0.3377
RFPR (sort by view count) 0.6632 0.5616 0.5226 0.5825 0.4456 0.5002 0.7607 0.5688 0.5757
RFPR (sort by selling count) 0.6248 0.5113 0.4672 0.5344 0.5029 0.5453 0.7682 0.6055 0.5700
RFPR (sort by manually-tuned formula) 0.6552 0.5421 0.4997 0.5657 0.4974 0.5422 0.7626 0.6007 0.5832
RFPR (sort by manually-tuned + IRF (Jung et al., 2007; Chakrabarti et al., 2008)) 0.6579 0.5437 0.5009 0.5675 0.4985 0.5431 0.7629 0.6015 0.5845
HPO (random; 500 samples) 0.7205 0.5817 0.5296 0.6106 0.4958 0.5392 0.7613 0.5988 0.6047
HPO (grid- extracted as ctr metrics best) 0.7326 0.5902 0.5367 0.6198 0.4999 0.5429 0.7649 0.6026 0.6112
HPO (grid- extracted as ctcvr metrics best) 0.7188 0.5817 0.5296 0.6100 0.5083 0.5503 0.7645 0.6077 0.6089
(Single Objective Tune)
MOHPER (TPE- ctr) 0.7312 0.5851 0.5300 0.6154 0.5029 0.5454 0.7598 0.6027 0.6091
MOHPER (TPE- ctcvr) 0.7293 0.5815 0.5261 0.6123 0.5166 0.5567 0.7646 0.6127 0.6125
(Separate Multi-Objective Tune)
MOHPER (TPE- ctr+ ctcvr) 0.7328 0.5875 0.5327 0.6177 0.5169 0.5575 0.7674 0.6139 0.6158
MOHPER (TPE- ctr+ ctcvr+ ctcar) 0.7315 0.5881 0.5342 0.6179 0.5184 0.5584 0.7668 0.6145 0.6162
MOHPER (TPE- ctr+ ctcvr+ ctcar; 500 samples) 0.7431 0.5950 0.5395 0.6259 0.5167 0.5570 0.7672 0.6136 0.6197
(Weighted Sum-Based Multi-Objective Tune)
MOHPER (TPE- ctr+ ctcvr) 0.7400 0.5941 0.5392 0.6244 0.5130 0.5539 0.7674 0.6114 0.6179
MOHPER (TPE- ctr+ ctcvr+ ctcar) 0.7228 0.5790 0.5251 0.6090 0.5232 0.5613 0.7678 0.6175 0.6132
MOHPER (TPE- ctr+ ctcvr+ ctcar; 500 samples) 0.7388 0.5926 0.5379 0.6231 0.5229 0.5618 0.7683 0.6176 0.6204
  • : Humans manually decide the weights of each model and choose CTR-based implicit relevance feedback (IRF) is added for ranking.
Table 3. Performance comparison of various models and optimization objectives on CPLP test data.
Method CTR nDCG@20 CTR nDCG@100 CTR prec@100 CTR avg metrics CTCVR nDCG@20 CTCVR nDCG@100 CTCVR prec@100 CTCVR avg metrics Overall avg metrics
RFPR (sort by view count) 0.6323 0.6673 0.7966 0.6987 0.5014 0.5400 0.8613 0.6342 0.6665
RFPR (sort by sell count) 0.5665 0.5471 0.5803 0.5646 0.5893 0.6201 0.8697 0.6930 0.6288
RFPR (sort by manually-tuned formula) 0.6263 0.6551 0.7786 0.6867 0.5514 0.5870 0.8672 0.6685 0.6776
HPO (random; 500 samples) 0.6340 0.6605 0.7806 0.6917 0.5680 0.6016 0.8692 0.6796 0.6857
HPO (grid- extracted as ctr metrics best) 0.6436 0.6697 0.7885 0.7006 0.5480 0.5831 0.8668 0.6660 0.6833
HPO (grid- extracted as ctcvr metrics best) 0.6305 0.6559 0.7767 0.6877 0.5679 0.6010 0.8688 0.6793 0.6835
(Single Objective Tune)
MOHPER (TPE- ctr) 0.6514 0.6764 0.7912 0.7063 0.5502 0.5835 0.8655 0.6664 0.6864
MOHPER (TPE- ctcvr) 0.6235 0.6522 0.7740 0.6832 0.5744 0.6058 0.8723 0.6842 0.6837
(Separate Multi-Objective Tune)
MOHPER (TPE- ctr+ ctcvr) 0.6321 0.6600 0.7800 0.6907 0.5738 0.6053 0.8722 0.6838 0.6872
MOHPER (TPE- ctr+ ctcvr+ ctcar) 0.6374 0.6634 0.7822 0.6943 0.5724 0.6056 0.8693 0.6824 0.6884
(Weighted Sum Based Multi-Objective Tune)
MOHPER (TPE- ctr+ ctcvr) 0.6320 0.6601 0.7800 0.6907 0.5760 0.6072 0.8724 0.6852 0.6880
MOHPER (TPE- ctr+ ctcvr+ ctcar) 0.6377 0.6656 0.7850 0.6961 0.5724 0.6028 0.8718 0.6823 0.6892
MOHPER (GP- ctr+ ctcvr+ ctcar) 0.6488 0.6712 0.7889 0.7013 0.5766 0.6073 0.8690 0.6843 0.6928
3.1.4. Samplers
In this study, we evaluate the performance of various hyperparameter optimization samplers to identify the most effective configuration for the proposed multi-objective optimization framework. The samplers under consideration include:
Random Sampler: This sampler employs a uniform random search strategy to explore the hyperparameter space without any prior knowledge, providing a baseline for comparison.
Grid Sampler: The grid search method systematically explores a predefined grid of hyperparameter values, ensuring exhaustive evaluation across a fixed set of possible configurations. While computationally expensive, it is useful for identifying optimal configurations within a constrained search space.
Gaussian Process (GP) Sampler: The GP sampler (Močkus, 1975) employs Gaussian processes (Rasmussen, 2003) to model the objective function and iteratively refine the search space by selecting hyperparameter configurations with high expected improvement. This method is particularly effective when function evaluations are costly and the search space is complex.
NSGA-III Sampler: The NSGA-III sampler (Deb and Jain, 2013) is an evolutionary algorithm designed for MOHPO. It applies a non-dominated sorting approach and maintains a diverse set of solutions to balance multiple objectives efficiently. This sampler is particularly known to be suited for problems with a large number of objectives and aims to achieve a well-distributed set of Pareto-optimal solutions.
Tree-structured Parzen Estimator (TPE) Sampler: The TPE sampler (Bergstra et al., 2011) uses a probabilistic model to guide the search, building a density estimate of the objective function and using it to propose new configurations. This model-based approach is designed to be more efficient than random or grid search, especially in high-dimensional spaces.
Each of these samplers is tested within the context of the multi-objective optimization problem to evaluate their efficacy in tuning hyperparameters for e-commerce retrieval tasks.
Table 4. Comparison of Sampler and Objective Type Performance for SERP and CPLP data.
Experiment Setups on SERP data on CPLP data
Sampler Objective def. type # Samples CTR avg. metrics CTCVR avg. metrics Overall avg. metrics CTR avg. metrics CTCVR avg. metrics Overall avg. metrics
Random - 500 0.6131 0.5982 0.6057 0.6956 0.6811 0.6884
Grid - # cases 0.6157 0.6078 0.6118 0.6954 0.6801 0.6877
NSGA-III Separately optimized 100 0.6130 0.6024 0.6077 0.6974 0.6774 0.6874
NSGA-III Separately optimized 500 0.6198 0.6166 0.6182 0.6988 0.6830 0.6909
NSGA-III Weighted sum-based 100 0.6117 0.6009 0.6063 0.6938 0.6810 0.6874
NSGA-III Weighted sum-based 500 0.6214 0.6167 0.6191 0.6919 0.6841 0.6880
GP Weighted sum-based 100 0.6140 0.6194 0.6167 0.7013 0.6843 0.6928
GP Weighted sum-based 500 0.6210 0.6197 0.6203 0.6952 0.6890 0.6921
TPE Separately optimized 100 0.6179 0.6145 0.6162 0.7010 0.6776 0.6893
TPE Separately optimized 500 0.6259 0.6136 0.6197 0.6942 0.6856 0.6899
TPE Weighted sum-based 100 0.6154 0.6124 0.6139 0.6961 0.6841 0.6901
TPE Weighted sum-based 500 0.6231 0.6176 0.6204 0.6911 0.6877 0.6894
3.1.5. Common Baselines on Offline Evaluation
For offline evaluation, we establish several baseline models commonly used in information retrieval tasks to provide comparative benchmarks for our Multi-Objective Hyperparameter Optimization (MOHPO) framework. These baselines are as follows:
BM25: A robust probabilistic retrieval model, BM25 (Robertson et al., 2009) ranks documents based on term frequency and inverse document frequency, making it a baseline for sparse keyword-based retrieval.
LLM-based Retrieval: This method leverages a large language model (LLM) to generate contextual embeddings, capturing semantic nuances in queries and item descriptions for dense retrieval. For Korean, we employ KoBERT [^7] https://github.com/SKTBrain/KoBERT, a fine-tuned variant of BERT (Devlin, 2018).
CLIP-based Retrieval: Utilizing a pretrained CLIP model (Radford et al., 2021), this method encodes queries and items into a shared embedding space. We implement two retrieval scenarios: query-to-image matching and query-to-item description matching. For Korean-language support, we adopt Jina-clip-v2 (Koukounas et al., 2024) and choose to apply 768 dims embeddings, since it does not harm the performance due to (Kusupati et al., 2022).
Hybrid Search: This approach integrates BM25’s sparse matching with vector-based semantic retrieval to enhance ranking performance by leveraging their complementary strengths. Similar hybrid approaches are explored in (Wang et al., 2021; Ghawi and Pfeffer, 2019).
Relevance Filtering with Popularity-based Ranking (RFPR): This baseline employs a two-stage ranking strategy. First, items are filtered based on query relevance (e.g., category or attribute matching). Then, the filtered set is ranked according to popularity, measured via historical user interaction metrics such as click-through rates, purchase frequencies, and cart additions. RFPR is widely used in e-commerce retrieval systems to prioritize high-engagement products, thereby improving overall relevance and user satisfaction. Additionally, we evaluate hybrid configurations that combine retrieval models with popularity-based ranking.
Each baseline method is applied to the evaluation dataset scenarios, as described in Section 3.1.1, allowing us to compare their effectiveness relative to our MOHPO framework.
3.1.6. Objectives and Metrics
As outlined in Section 2.2, our primary optimization goals are CTCVR and CTR, with auxiliary objectives aimed at capturing both user engagement and relevance between products and items, as well as actual conversion intent. To assess the effectiveness of the Multi-Objective Hyperparameter Optimization (MOHPO) framework, we compare it with baseline single-objective models and various multi-objective configurations under different sampler settings. Offline evaluations are conducted on holdout test datasets, while online A/B testing measures real-time performance within a production environment. For offline evaluations, while monitoring multiple metrics that correspond to training objectives, we mainly present nDCG@20, nDCG@100, and precision@100 to succinctly report results. These metrics serve as reliable indicators of user experience on our OHouse platform, as they closely align with engagement behaviors and satisfaction.
3.2. Experimental Results and Discussion
3.2.1. RQ1a. Offline Performance Comparison on Balancing Multi-objective E-commerce Metrics
To assess the efficacy of the proposed MOHPER framework, we benchmark its performance against several baseline approaches. Specifically, we examine its application in two distinct e-commerce scenarios: retrieval with keyword queries for search engine result pages (SERP) and retrieval based on category IDs for product list pages (CPLP). The experimental setup involves optimizing three primary metrics (nDCG@20, precision@100), and varying relevance targets across the following five multi-objective scenarios: (a) CTR (2 objectives), (b) CTCVR (2 objectives), (c) CTR+ CTCVR (4 objectives), (d) CTR+ CTCVR+ CTCAR (6 objectives). For multi-objective training, we adopt two approaches: (i) separate multi-objective tuning, where each objective is trained individually within a multi-objective framework, and (ii) weighted sum-based multi-objective tuning, which aggregates metrics into a single weighted objective for optimization. To ensure fairness in objective selection, we limit hyperparameter optimization (HPO) methods to 100 samples by default. For a broader comparison, we also evaluate further sampled configurations. Grid search baselines explore 1,024 configurations for SERP and 512 for CPLP.
The comparative results, presented in Table 2 and Table 3, demonstrate that MOHPER consistently outperforms baseline methods such as RFPR, random search, and grid search across all configurations. While single-objective approaches (a and b) achieve strong performance for their respective primary objectives (e.g., CTR in scenario a), they fail to optimize complementary metrics effectively (e.g., CTCVR in scenario a). In contrast, multi-objective configurations, particularly scenario (c), exhibit a more balanced optimization across CTR and CTCVR, yielding superior overall performance. Furthermore, the inclusion of auxiliary objectives in scenarios (d) and (e) enhances the robustness of the results, enabling more generalized performance across diverse evaluation metrics. These findings highlight the effectiveness of MOHPER in addressing complex multi-objective optimization problems in e-commerce retrieval systems.
3.2.2. RQ1b. Performance Difference According to Samplers and Setups
To evaluate the effectiveness of different samplers, we compared their performance across varying sample sizes to determine their suitability for specific optimization scenarios. The experimental setup involved training the same objectives outlined in 3.2.1, while varying the type of samplers and the number of samples. The comparative results are summarized in Table 4. Among the evaluated samplers, the TPE sampler consistently demonstrated superior performance across all cases and proved to be applicable in diverse scenarios. In contrast, the GP sampler excelled in settings with a small number of samples; however, it is inherently limited to single-objective optimization unless multi-objective problems are reformulated into a single objective using weighted summation. The NSGA-III sampler, while applicable to multi-objective scenarios without such reformulation, exhibited reduced effectiveness when the number of samples was constrained. The results indicate that increasing the number of samples generally enhances the performance of samplers, albeit at the cost of higher computational overhead. To achieve a more efficient trade-off between performance and computational cost, we introduced a cumulative training strategy, as detailed in the experiments of 3.2.4. This approach leverages incremental sampling and training to boost performance while maintaining computational efficiency.
3.2.3. RQ2. Hybrid Retrieval Arbitration
Modern e-commerce systems employ complex retrieval strategies that involve a wide range of feature selection choices, encompassing both sparse (lexical) and dense (vector-based) embedding retrieval methods and item popularity. To illustrate the interplay among three core feature types in e-commerce retrieval, sparse lexical similarity, dense semantic similarity, and item popularity, we utilize BM25, KoBERT, and business quantity features of items, respectively. BM25 represents sparse lexical similarity, while LLM captures dense semantic similarity. Additionally, business quantity features highlight item popularity, a critical aspect in real-world e-commerce retrieval. To underscore the significance of popularity-based features over similarity-based features in production systems, we also include RFPR baselines.
The evaluation metrics, nDCG@20 and Recall@100, are calculated on query-item pairs using CTR or CTCVR as relevance labels. The results, summarized in Table 5, demonstrate that the proposed MOHPER framework effectively balances the contributions of sparse and dense similarities with popularity features. This adaptability enables MOHPER to align retrieval strategies with user preferences and business objectives, optimizing performance across diverse e-commerce scenarios.
Table 5. Performance comparison for retrieval strategies on SERP data
Method CTR nDCG@20 CTCVR nDCG@20 CTR recall@100 CTCVR recall@100 Overall Avg.
BM25 0.2270 0.0831 0.2516 0.3951 0.2392
ANN - (w/ KoBERT) 0.3753 0.3227 0.2270 0.4604 0.3463
ANN + BM25 - (grid search) 0.3855 0.3502 0.2139 0.4551 0.3512
RFPR - (by selling) 0.6248 0.5029 0.4108 0.7682 0.5767
ANN + RFPR 0.6253 0.5026 0.4125 0.7681 0.5771
MOHPER 0.7400 0.5178 0.4718 0.7685 0.6245
3.2.4. RQ3. Effectiveness of Voting and Cumulative Learning
Table 6. Performance comparison across cumulative training stage with ablation analysis of voting in the meta-stage
Separate Objectives Tune Weighted Sum-based
Cumulative Stage (# sample/# total) (default: w/ voting) on SERP avg. metrics on CPLP avg. metrics on SERP weighted score on CPLP weighted score
[Stage 0; 500/500] (wo/ voting) 0.6221 0.6759 0.9365 1.0884
[Stage 0; 500/500] 0.6257 0.6778 0.9374 1.0884
[Stage 0; 100/100] 0.6223 0.6756 0.9344 1.0762
[Stage 1; 100/200] 0.6249 0.6756 0.9344 1.0920
[Stage 2; 100/300] 0.6257 0.6786 0.9354 1.0920
[Stage 3; 100/400] (wo/ voting) 0.6244 0.6781 0.9354 1.0864
[Stage 3; 100/400] 0.6257 0.6786 0.9370 1.0920
  • : No incremental performance is observed from the previous stage, likely due to stochastic variations in sampling logic.
This section demonstrates the effectiveness of voting and cumulative learning strategies in optimization. The results, presented in 6, indicate that voting mechanisms can improve performance compared to no-voting naive optimization. This approach, consistent with the benefits observed in ensemble methods (Chan and Stolfo, 1995; Bauer and Kohavi, 1999; Dietterich, 2000), reduces variance and enhances generalization. On the other hand, cumulative learning further improves performance by iteratively building upon prior knowledge, which reduces the need for large sample sizes and minimizes redundant noise introduced by samplers. We hypothesize that this approach effectively mitigates redundant information, particularly in samplers like Tree-structured Parzen Estimator (TPE), which does not explicitly account for variance during its optimization process. As training progresses, the average differentiation between high-performing (good) and low-performing (bad) samples diminishes due to convergence. Cumulative learning counters this effect by selectively retaining high-quality samples and filtering out redundant or low-quality ones, thereby restoring differentiation and enabling more efficient exploration of the hyperparameter space. Specifically, for TPE, this refined focus on promising regions enhances its ability to identify optimal configurations, leading to uplift performance. Moreover, an additional advantage of splitting # samples into a series of mini-cumulative stages is improved debuggability, and the simplicity of implementation between stages further enhances the practicality and effectiveness of this approach.
3.2.5. RQ4. Online AB Test
We present empirical results from applying our proposed methods to various e-commerce experiments on OHouse retrieval tasks. The experiments on SERP (Search Engine Results Page) were conducted over two weeks, while those on CPLP (Category Product List Page) spanned three weeks. To isolate the performance of MOHPER, we maintained the original search space from previous setups and focused solely on adjusting service hyperparameters. These adjustments included modulating model/feature weights and refining model/feature selections for retrieval or ranking. Table 7. summarizes the results, focusing on key performance metrics: CTR, CTCAR, CTCVR, and per-query engagement rates (Click/Q, Cart/Q, and Purchase/Q). While CTR on CPLP showed a slight decline, all other metrics exhibited consistent improvements, demonstrating MOHPER’s effectiveness in optimizing retrieval performance across multiple objectives. Notably, these gains were achieved without increasing computational costs, as our approach modified only hyperparameters without introducing additional features. Although we do not provide additional online empirical results for fairness and clarification, we note that subsequent series of experiments expanding the search space were successfully launched following these conditioned successes.
Table 7. Online empirical results of applying MOHPER to SERP and CPLP on OHouse retrieval tasks, meaning ${}^{*}:(p<0.1),^{**}:(p<0.05),$ and ${}^{***}:(p<0.01)$
Exp. CTR gain(%) CTCAR gain(%) CTCVR gain(%) Click/Q gain(%) Cart/Q gain(%) Purchase/Q gain(%)
on SERP 0.95%*** 0.75%*** 0.30% 1.23%*** 0.98%* 0.75%*
on CPLP -0.51%*** 1.16%*** 1.10%* -0.22%** 1.45%** 1.62%*
4. Conclusions
This paper presents MOHPER, a novel multi-objective hyperparameter optimization (HPO) framework tailored to the demands of e-commerce retrieval, specifically optimizing for critical performance metrics such as CTR, CTCAR, and CTCVR, which are pivotal for enhancing user engagement and conversions in e-commerce systems. As the first multi-objective HPO framework applied to retrieval and ranking tasks in this domain, MOHPER effectively tackles the inherent complexity of concurrently optimizing diverse and competing metrics in e-commerce. Through extensive evaluation, we demonstrated MOHPER’s ability to achieve superior stability, convergence, and robustness in multi-objectives optimization across various HPO sampling strategies, achieving significant improvements in CTR, CTCVR, and related objectives (RQ1). The integration of hybrid optimization mechanisms further enhanced MOHPER’s performance, delivering measurable gains in relevance and user engagement, thus advancing multi-objective optimization for e-commerce (RQ2). Additionally, we developed advanced techniques such as meta-configuration voting and cumulative training, which uplift MOHPER’s adaptability by shortening sampling costs (RQ3) and also provide inspection changes between each cumulative stage. Lastly, large-scale A/B testing conducted on OHouse’s dataset demonstrated MOHPER’s scalability and validated its practical utility, solidifying its role as a foundational optimization tool within the platform (RQ4). In future work, extending MOHPER’s capabilities could involve exploring new search spaces for training diverse machine learning algorithms and refining the design of objective functions to incorporate factors such as the freshness and diversification of retrieved items.